# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v2

@@ 8 Conclusion
We have shown that recognizing regular languages and graph connectivity, two key problems inexpressible by fixed-depth transformers, become expressible if the depth of the transformer can grow very slightly (logarithmically) with the context length by repeating layers. This implies transformers with fixed depth $d$ can solve these problems up to bounded context lengths of $2^{O(d)}$. Further, we showed that scaling depth to solve these problems is more efficient than scaling width (which requires superpolynomial increase) or scaling chain-of-thought steps (which requires superlogarithmic increase). As dynamic test-time compute methods have become popular for building more powerful reasoning models such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), it would be interesting to explore whether universal transformers can realize this theoretical efficiency to provide more efficient long-context reasoning than chain-of-thought steps in practice.
While growing depth enables transformers to solve some key problems outside $\mathsf{TC}^{0}$, there are limitations on the types of problems log depth should enable solving. Unless $\mathsf{NC}=\mathsf{P}$, log-depth (or even polylog-depth) transformers cannot express $\mathsf{P}$-complete problems including solving linear equalities, in-context context-free language recognition, circuit evaluation, and determining the satisfiability of Horn clauses (Greenlaw et al., 1991). In future work, it would interesting to study the depth required for transformers to solve other interesting problems in $\mathsf{NC}$ including context-free recognition (generalizing regular languages; Theorem 1), which is in $\mathsf{AC}^{1}$ (Ruzzo, 1981; Venkateswaran, 1991) and boolean formula evaluation, which is $\mathsf{NC}^{1}$-complete (Buss, 1987). This would help us better understand the degree to which repeating layers can be used as a form of interence-time compute.
@@ A.1 Residual Stream Storage Interface
In order to retrieve a value $z$ with masked pre-norm (rather than just its sign), we can instead represent $z$ as a 4-dimensional vector $\psi(z)=\langle z,1,-z,-1\rangle$. Then, pre-norm masked to only this vector will return $\phi(z)=\psi(z)/\sqrt{z^{2}+1}$. Scalars $z$ stored as $\psi(z)$ or $\phi(z)$ in the residual stream can be trivially retrieved as $\phi(z)$ by masked pre-norm:
@@
The first layer computes $z$ and stores $\psi(z)$ at blank indices $I$ in the residual stream, producing $\mathbf{h}^{\prime}$. Thus, the second layer can read $\phi(z)$ with masked pre-norm via Lemma 2 and can also recompute $z$ from $\mathbf{h}$, which is a subspace of $\mathbf{h}^{\prime}$. At this point, it outputs $-\psi(z)+\phi(z)$ at indices $I$, which leads to $\mathbf{h}^{\prime\prime}$ storing $\phi(z)$ at $I$. ∎
@@ A.2 Clearing Stored Values
In the repeated layers of a universal transformer, we will need to overwrite the values stored at particular indices in the residual stream. That is, if $[\mathbf{h}]_{I}=\mathbf{x}$, it will be useful to produce $\mathbf{h}^{\prime}$ such that $[\mathbf{h}^{\prime}]_{I}=\mathbf{y}$ instead. The following lemmas will help implement constructions of this form.
@@
If a unit-norm vector $\phi$ is stored in $\mathbf{h}$ at $I$, there exists a feedforward sub-layer that removes $\phi$, i.e., produces $\mathbf{h}^{\prime}$ such that $[\mathbf{h}^{\prime}]_{i}=\vec{0}$.
@@
The layer reads $\phi$ via masked pre-norm and writes $-\phi$ to $\mathbf{h}$ at $I$, setting $[\mathbf{h}]_{I}=\phi-\phi=\vec{0}$. ∎
Combining Lemma 5 with a parallel layer that stores some new value at $I$, we see that we can effectively overwrite values at $I$ rather than just deleting them.
@@
The second layer is a copy of the initial layer that considers the subvector $\mathbf{h}$ of $\mathbf{h}^{\prime}$ as its input and where all signs are flipped. Thus, it outputs $-\delta$, which guarantees that the final residual stream is $\mathbf{h}^{\prime\prime}=\mathbf{h}+\delta-\delta=\mathbf{h}$. ∎
@@ A.3 Computing Position Offsets
Assume a transformer stores $\mathbbm{1}[i=0]$ and $\mathbbm{1}[i<k]$ in the residual stream. Then, with 1 layer, it possible to add $\phi(i-k)$ in the residual stream at indices $i\geq k$.
@@
We construct two attention heads. The first is uniform with value $\mathbbm{1}[j=0]$, and thus computes $1/i$. The second is uniform with value $\mathbbm{1}[j\geq k]$, and thus computes $(i-k)/i$. We then use a feedforward layer to compute $\phi((i-k)/i,1/i)=\phi(i-k)$ and store it in the residual stream. ∎
The precondition that we can identify the initial token [cf. Merrill and Sabharwal, 2024] is easy to meet with any natural representation of position, including $1/i$ or $\phi(i)$, as we can simply compare the position representation against some constant.
We assume that the positional encodings used by the model allow detecting the initial token [Merrill and Sabharwal, 2024]. One way to enable this would simply be to add a beginning-of-sequence token, although most position embeddings should also enable it directly.
@@ A.4 Equality Checks
After computing $x,y$ in a self-attention layer, we write $x-y$ to a temporary cell in the residual stream. The feedforward sublayer reads $\sigma_{1}=\mathrm{sgn}(x-y)$, computes $z=1-\mathrm{ReLU}(\sigma_{1})-\mathrm{ReLU}(-\sigma_{1})$, and writes $z$ to the residual stream.
@@
The next transformer layer then recomputes $y-x$ and adds it to the intermediate memory cell, which sets it back to 0. Thus, the output is correct and intermediate memory is cleared. ∎
Appendix B Division Construction Correctness
@@
The proof of Lemma 1 presents the full construction to implement division in a transformer. For space, we omitted a full proof of correctness for the construction, which we now present.
Proof of Correctness.
@@
In the first layer, suppose first that $i$ is a multiple of $m$. In this case, there exists a position $j^{*}\leq i$ such that $i=mj^{*}$, which means the query $q_{i}=\phi(i/m)=\phi(j^{*})$ exactly matches the key $k_{j^{*}}$. The head will thus return $v_{j^{*}}=\phi(j^{*})=\phi(i/m)$, representing precisely the quotient $i/m$. Further, the equality check will pass, making $e_{i}=1$. The layer thus behaves as intended when $i$ is a multiple of $m$. On the other hand, when $i$ is not a multiple of $m$, no such $j^{*}$ exists. The head will instead attend to some $j$ for which $i\neq mj$ and therefore $\phi(i/m)\neq\phi(j)$, making the subsequent equality check fail and setting $e_{i}=0$, as intended.
@@
For the second layer, let $h^{2}_{i}$ denote the head’s output at position $i$. By construction, $q_{i}\cdot k_{j}=e_{j}-[\phi(j)]_{0}$ where $[\phi(j)]_{0}=j/\sqrt{2j^{2}+2}$ is the first coordinate of $\phi(j)$. Note that $[\phi(j)]_{0}\in[0,1)$ for positions $j\leq i$ and that it is monotonically increasing in $j$. It follows that the dot product is maximized at the largest $j\leq i$ such that $e_{j}=1$, i.e., at the largest $j\leq i$ that is a multiple of $m$. This $j$ has the property that $\lfloor i/m\rfloor=j/m$. Thus, the head at this layer attends solely to this $j$. Recall that the value $v_{j}$ at this position is $h^{1}_{j}=\phi(j/m)$, which thus equals $\phi(\lfloor i/m\rfloor)$. The head’s output, $h^{2}_{i}$, is therefore $\phi(\lfloor i/m\rfloor)$, as intended, which is stored in the residual stream.
The correctness of the third and fourth layer is easy to verify.
@@
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$; each $b^{k}_{j}$ for $k\in\{1,2\}$ can be retrieved from the residual stream. The query-key product achieves its upper bound of 2 exactly when two conditions hold: $\lfloor i/m\rfloor=\lfloor j/m\rfloor$ and $b^{1}_{j}=1$. Thus, the head attends from $i$ to all $j\leq i$ that store the same quotient as $i$ and also have $b^{1}_{j}=1$. To make this clearer, let’s write $i$ as $i=b^{\prime}m+c^{\prime}$ for some $c^{\prime}<m$. The query-key dot product is then maximized precisely at the $c^{\prime}$ positions $j$ in $\{b^{\prime}m+1,b^{\prime}m+2,\ldots,b^{\prime}m+c^{\prime}\}$, for all of which $\lfloor j/m\rfloor=\lfloor i/m\rfloor=b^{\prime}$; note that $b^{\prime}m$ is not included in this list as $b^{1}_{j}=0$ when $j=b^{\prime}m$. Of these positions, only $b^{\prime}m+1$ has the property that the quotient there is not the same as the quotient two position earlier, as captured by the value $v_{j}=1-b^{2}_{j}$. Thus, the value $v_{j}$ is $1$ among these positions only at $j=b^{\prime}m+1$, and 0 elsewhere.
@@
On the other hand, if $m$ does divide $i$ (which can be checked with a separate, parallel head), we write $\psi(0)$ to the residual stream, which, when read by the next layer, will yield $\phi(0)=\phi(i\bmod m)$.
The sixth layer attends with query $q_{i}=\phi(a_{i})$, key $k_{j}=\phi(j)$, and value $v_{j}=\langle\phi(\lfloor j/m\rfloor),\phi(j\bmod m)\rangle$, where both components of the value have been previously computed and stored in the residual stream in layers two and five. Since $a_{i}\leq i$, the query matches the key at exactly one position $j=a_{i}$, and the head retrieves $\langle\phi(\lfloor a_{i}/m\rfloor),\phi(a_{i}\bmod m)\rangle$, which is precisely $\langle\phi(b_{i}),\phi(c_{i})\rangle$. The layer can thus store $\phi(b_{i})$ and $\phi(c_{i})$ to the residual stream at position $i$.
That the seventh “cleanup” layer operates as desired is easy to see from the construction. ∎
@@ Appendix C Regular Language Recognition Proof
Regular language recognition can be framed as multiplying a sequence of elements in the automaton’s transition monoid [Myhill, 1957, Thérien, 1981]. It thus suffices to show how $n$ elements in a finite monoid can be multiplied with $\Theta(\log n)$ depth. We show how a log-depth universal transformer can implement the standard binary tree construction [Barrington and Thérien, 1988, Liu et al., 2023, Merrill et al., 2024] where each level multiplies two items, meaning the overall depth is $\Theta(\log\lvert w\rvert)$. We will represent a tree over the input tokens within the transformer. Each level of the tree will take 8 transformer layers. We define a notion of active tokens: at level 0, all tokens are active, and, at level $\ell$, tokens at $t\cdot 2^{\ell}-1$ for any $t$ will remain active, and all other tokens will be marked as inactive. As an invariant, active token $i=t\cdot 2^{\ell}-1$ in level $\ell$ will store a unit-norm vector $\delta^{\ell}_{i}$ that represents the cumulative product of tokens from $i-2^{\ell}+1$ to $i$.
@@
We now proceed by induction over $\ell$, defining the behavior of non-\$ tokens at layers that make up level $\ell$. The current group element $\delta^{\ell}_{i}$ stored at active token $i$ is, by inductive assumption, the cumulative product from $i-2^{\ell}+1$ to $i$. Let $\alpha^{\ell}_{i}$ denote that token $i$ is active. By Lemma 7 we use a layer to store $i-1$ at token $i$. The next layer attends with query $\phi(i-1)$, key $\phi(j)$, and value $\delta^{\ell}_{j}$ to retrieve $\delta^{\ell}_{i-1}$, the group element stored at the previous token. Finally, another layer attends with query $\vec{1}$, key $\langle\phi(j)_{1},\alpha^{\ell}_{i}\rangle$, and value $\delta^{\ell}_{j-1}$ to retrieve the group element $\delta^{\ell}_{j^{*}}$ stored at the previous active token, which represents the cumulative product from $i-2\cdot 2^{\ell}+1$ to $i-2^{\ell}$. Next, we will use two layers to update $\delta^{\ell}_{i}\leftarrow\delta^{\ell+1}_{i}$ and $\delta^{\ell}_{j}\leftarrow\vec{0}$, which is achieved as follows. First, we assert there exists a single feedforward layer that uses a table lookup to compute $\delta^{\ell}_{j^{*}},\delta^{\ell}_{i}\mapsto d$ such that $d/\lVert d\rVert=\delta^{\ell}_{j^{*}}\cdot\delta^{\ell}_{i}=\delta^{\ell+1}_{i}$. Next, we invoke Lemma 6 to construct a layer that adds $d$ to an empty cell of the residual stream and then another layer that deletes it. This second layer can now read both $\delta^{\ell}_{i},\delta^{\ell}_{j^{*}}$ and $\delta^{\ell+1}_{i}$ (from $d$) as input, and we modify it to add $\delta^{\ell+1}_{i}-\delta^{\ell}_{i}$ to $\delta^{\ell}_{i}$, changing its value to $\delta^{\ell+1}_{i}$. Similarly, we modify it to add $-\delta^{\ell}_{j^{*}}$ to $\delta^{\ell}_{j^{*}}$ to set it to 0. A feedforward network then subtracts $\delta^{\ell}_{i}$ from the residual stream and adds $\delta^{\ell}_{i}\cdot\delta^{\ell}_{j}$. This requires at most 4 layers.
@@
To determine activeness in layer $\ell+1$, each token $i$ attends to its left to compute $c_{i}/i$, where $c_{i}$ is the prefix count of active tokens, inclusive of the current token. We then compute $\phi(c_{i}/i,1/i)=\phi(c_{i})$ and store $c_{i}$ it temporarily in the residual stream. At this point, we use Lemma 1 to construct 7 layers that compute $c_{i}\bmod 2$ with no storage overhead. The current token is marked as active in layer $\ell+1$ iff $c_{i}=0\mod 2$, which is equivalent to checking whether $i=t\cdot 2^{\ell}-1$ for some $t$. In addition to updating the new activeness $\alpha^{\ell+1}_{i}$, we also persist store the previous activeness $\alpha^{\ell}_{i}$ in a separate cell of the residual stream and clear $c_{i}$. This requires at most 8 layers.
Finally, we describe how to aggregate the cumulative product at the \$currency-dollar\\$\$ token, which happens in parallel to the behavior at other tokens. Let $\delta^{\ell}_{\$}$ be a monoid element stored at \$currency-dollar\\$\$ that is initialized to the identity and will be updated at each layer. Using the previously stored value $i-1$, we can use a layer to compute and store $\alpha^{\ell}_{i-1}$ and $\alpha^{\ell+1}_{i-1}$ at each $i$. A head then attends with query $\vec{1}$, key $\langle\phi(j)_{1},10\cdot\alpha^{\ell}_{i-1}\rangle$, and value $\langle(1-\alpha^{\ell+1}_{j-1})\cdot\delta^{\ell+1}_{j-1}\rangle$. This retrieves a value from the previous active token $j$ at level $\ell$ that is $\delta^{\ell}_{j}$ if $j$ will become inactive at $\ell+1$ and $\vec{0}$ otherwise. Iff $\delta^{\ell}_{j}$ is retrieved, a feedforward network subtracts $\delta^{\ell}_{\$}$ from the residual stream and adds $\delta^{\ell}_{j}\cdot\delta^{\ell}_{\$}$. This guarantees that whenever a tree is deactivated, its cumulative product is incorporated into $\delta^{\ell}_{\$}$. Thus, after $\ell=\lceil\log_{2}\lvert w\rvert\rceil+1$ levels, $\delta^{\ell}_{\$}$ will be the transition monoid element for $w$. We can use one additional layer to check whether this monoid element maps the initial state to an accepting state using a finite lookup table. Overall, this can be expressed with 8 layers repeated $\lceil\log_{2}\lvert w\rvert\rceil$ times and 9 final layers (to implement the additional step beyond $\lceil\log n\rceil$). ∎

### v3

@@ 8 Conclusion
We have shown that recognizing regular languages and graph connectivity, two key problems inexpressible by fixed-depth transformers, become expressible if the depth of the transformer can grow very slightly (logarithmically) with the context length by repeating layers. This implies transformers with fixed depth $d$ can solve these problems up to bounded context lengths of $2^{O(d)}$. Further, we showed that scaling depth to solve these problems is more efficient than scaling width (which requires superpolynomial increase) or scaling chain-of-thought steps (which requires superlogarithmic increase). As dynamic test-time compute methods have become popular for building more powerful reasoning models such as OpenAI o1 (o1-short) and DeepSeek-R1 (deepseekr1-short), it would be interesting to explore whether universal transformers can realize this theoretical efficiency to provide more efficient long-context reasoning than chain-of-thought steps in practice.
While growing depth enables transformers to solve some key problems outside $\mathsf{TC}^{0}$, there are limitations on the types of problems log depth should enable solving. Unless $\mathsf{NC}=\mathsf{P}$, log-depth (or even polylog-depth) transformers cannot express $\mathsf{P}$-complete problems including solving linear equalities, in-context context-free language recognition, circuit evaluation, and determining the satisfiability of Horn clauses (greenlaw1991compendium). In future work, it would interesting to study the depth required for transformers to solve other interesting problems in $\mathsf{NC}$ including context-free recognition (generalizing regular languages; Theorem˜1), which is in $\mathsf{AC}^{1}$ (ruzzo-1981-uniform; venkateswaran-1991-properties) and boolean formula evaluation, which is $\mathsf{NC}^{1}$-complete (buss-1987-bfvp). This would help us better understand the degree to which repeating layers can be used as a form of interence-time compute.
Limitations
We have given constructions whereby looped transformers can express $\mathsf{NC}^{1}$-hard problems, though we have not considered looped transformers’ inductive biases and learning dynamics, which are also important in practice beyond expressivity. saunshi2025reasoning empirically study the inductive biases of looped transformers, suggesting looped transformers may generalize in ways favorable for reasoning tasks. Our experiments in Section˜7 were conducted with non-looped transformers. It would be interesting in future work to evaluate how the depth requirements change if we enforce that the transformer must re-use the same weights across layers.
@@ A.1 Residual Stream Storage Interface
In order to retrieve a value $z$ with masked pre-norm (rather than just its sign), we can instead represent $z$ as a 4-dimensional vector $\psi(z)=\langle z,1,-z,-1\rangle$. Then, pre-norm masked to only this vector will return $\phi(z)=1/\sqrt{2}\cdot\psi(z)/\sqrt{z^{2}+1}$. Scalars $z$ stored as $\psi(z)$ or $\phi(z)$ in the residual stream can be trivially retrieved as $\phi(z)$ by masked pre-norm:
@@
The first layer computes $z$ and stores $\psi(z)$ at blank indices $I$ in the residual stream, producing $\mathbf{h}^{\prime}$. Thus, the second layer can read $\phi(z)$ with masked pre-norm via Lemma˜2 and can also recompute $z$ from $\mathbf{h}$, which is a subspace of $\mathbf{h}^{\prime}$. At this point, it outputs $\psi(z)+\phi(z)$ at indices $I$, which leads to $\mathbf{h}^{\prime\prime}$ storing $\phi(z)$ at $I$. ∎
@@ A.2 Clearing Stored Values
In the repeated layers of a universal transformer, we will need to overwrite the values stored at particular indices in the residual stream. That is, if $\mathbf{h}]_{I}=\mathbf{x}$, it will be useful to produce $\mathbf{h}^{\prime}$ such that $\mathbf{h}^{\prime}]_{I}=\mathbf{y}$ instead. The following lemmas will help implement constructions of this form.
@@
If a unit-norm vector $\phi$ is stored in $\mathbf{h}$ at $I$, there exists a feedforward sub-layer that removes $\phi$, i.e., produces $\mathbf{h}^{\prime}$ such that $\mathbf{h}^{\prime}]_{i}=\vec{0}$.
@@
The layer reads $\phi$ via masked pre-norm and writes $\phi$ to $\mathbf{h}$ at $I$, setting $\mathbf{h}]_{I}=\phi-\phi=\vec{0}$. ∎
Combining Lemma˜5 with a parallel layer that stores some new value at $I$, we see that we can effectively overwrite values at $I$ rather than just deleting them.
@@
The second layer is a copy of the initial layer that considers the subvector $\mathbf{h}$ of $\mathbf{h}^{\prime}$ as its input and where all signs are flipped. Thus, it outputs $\delta$, which guarantees that the final residual stream is $\mathbf{h}^{\prime\prime}=\mathbf{h}+\delta-\delta=\mathbf{h}$. ∎
@@ A.3 Computing Position Offsets
For a fixed value $k$, assume that at each position $i$, a transformer stores $\mathbbm{1}[i=0]$ and $\mathbbm{1}[i<k]$ in the residual stream. Then, there exists a layer that adds $\phi(i-k)$ to the residual stream at token indices $i\geq k$.
@@
We construct two attention heads. The first head is uniform with value $j$ as $\mathbbm{1}[j=0]$. Thus, the head computes $h_{1}=1/i$. The second head is uniform with value $j$ as $\mathbbm{1}[j\geq k]$, and thus computes $h_{2}=(i-k)/i$. We then use a feedforward layer to compute:
$\phi(h_{1},h_{2})=\phi(\frac{i-k}{i},\frac{1}{i})=\phi(i-k).$
The resulting value is then written to the residual stream. ∎
The precondition that we can identify the initial token [cf. merrill-sabharwal-2024-cot] is easy to meet with any natural representation of position, including $1/i$ or $\phi(i)$, as we can simply compare the position representation against some constant.
We assume that the positional encodings used by the model allow detecting the initial token [merrill-sabharwal-2024-cot]. One way to enable this would simply be to add a beginning-of-sequence token, although most position embeddings should also enable it directly.
@@ A.4 Equality Checks
After computing $x,y$ in a self-attention layer, we write xyx-y to a temporary cell in the residual stream. The feedforward sublayer reads $\sigma_{1}=\mathrm{sgn}(x-y)$, computes $z=1-\mathrm{ReLU}(\sigma_{1})-\mathrm{ReLU}(-\sigma_{1})$, and writes $z$ to the residual stream.
@@
The next transformer layer then recomputes yxy-x and adds it to the intermediate memory cell, which sets it back to 0. Thus, the output is correct and intermediate memory is cleared. ∎
Appendix B Correctness of Division Construction Attention Heads
@@
The proof of Lemma˜1 presents the full construction to implement division in a transformer. For space, we omitted a full proofs of correctness for attention heads in the construction, which we now present. We expect some of these techniques could be reused, though we have not stated them as generally as the gadgets in Appendix˜A.
Lemma 9 (First Layer of Lemma˜1).
Let $h^{1}_{i}$ be an attention head computed with query $q_{i}=\phi(i/m)$ and keys/values $k_{j}=v_{j}=\phi(j)$. Let $e_{i}=\mathbb{I}(h_{i}^{1}=\phi(i/m))$. Then $e_{i}=1$ if and only if $i$ is a multiple of $m$. Furthermore, if $e_{i}=1$, then $h^{1}_{i}=\phi(i/m)$.
Proof.
@@
Suppose first that $i$ is a multiple of $m$. In this case, there exists a position $j^{*}\leq i$ such that $i=mj^{*}$, which means the query $q_{i}=\phi(i/m)=\phi(j^{*})$ exactly matches the key $k_{j^{*}}$. The head will thus return $v_{j^{*}}=\phi(j^{*})=\phi(i/m)$, representing precisely the quotient $i/m$. Further, the equality check will pass, making $e_{i}=1$. The layer thus behaves as intended when $i$ is a multiple of $m$.
On the other hand, when $i$ is not a multiple of $m$, no such $j^{*}$ exists. The head will instead attend to some $j$ for which $i\neq mj$ and therefore $\phi(i/m)\neq\phi(j)$, making the subsequent equality check fail and setting $e_{i}=0$, as intended. ∎
Lemma 10 (Second Layer of Lemma˜1).
Let $h^{2}_{i}$ be an attention head computed with $q_{i}=\langle 1,1\rangle$, key $k_{j}=\langle e_{j},[\phi(j)]_{0}\rangle$, and value $v_{j}=h^{1}_{j}$. Then $h^{2}_{i}=\phi(\lfloor i/m\rfloor)$ .
Proof.
@@
By construction, $q_{i}\cdot k_{j}=e_{j}-[\phi(j)]_{0}$ where $\phi(j)]_{0}=j/\sqrt{2j^{2}+2}$ is the first coordinate of $\phi(j)$. Note that $\phi(j)]_{0}\in[0,1)$ for positions $j\leq i$ and that it is monotonically increasing in $j$. It follows that the dot product is maximized at the largest $j\leq i$ such that $e_{j}=1$, i.e., at the largest $j\leq i$ that is a multiple of $m$. This $j$ has the property that $\lfloor i/m\rfloor=j/m$. Thus, the head at this layer attends solely to this $j$. Recall that the value $v_{j}$ at this position is $h^{1}_{j}=\phi(j/m)=\phi(\lfloor i/m\rfloor)$. ∎
Lemma 11 (Fifth Layer of Lemma˜1).
Let $h^{5}_{i}$ be an attention head computed with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$. Then $h^{5}_{i}=\phi(i\bmod m)$.
Proof.
@@
The query-key product achieves its upper bound of 2 exactly when two conditions hold: $\lfloor i/m\rfloor=\lfloor j/m\rfloor$ and $b^{1}_{j}=1$. Thus, the head attends from $i$ to all $j\leq i$ that store the same quotient as $i$ and also have $b^{1}_{j}=1$. To make this clearer, let’s write $i$ as $i=b^{\prime}m+c^{\prime}$ for some $c^{\prime}<m$. The query-key dot product is then maximized precisely at the $c^{\prime}$ positions $j$ in $\{b^{\prime}m+1,b^{\prime}m+2,\ldots,b^{\prime}m+c^{\prime}\}$, for all of which $\lfloor j/m\rfloor=\lfloor i/m\rfloor=b^{\prime}$; note that $b^{\prime}m$ is not included in this list as $b^{1}_{j}=0$ when $j=b^{\prime}m$. Of these positions, only $b^{\prime}m+1$ has the property that the quotient there is not the same as the quotient two position earlier, as captured by the value $v_{j}=1-b^{2}_{j}$. Thus, the value $v_{j}$ is $1$ among these positions only at $j=b^{\prime}m+1$, and 0 elsewhere.
@@
On the other hand, if $m$ does divide $i$ (which can be checked with a separate, parallel head), we write $\psi(0)$ to the residual stream, which, when read by the next layer, will yield $\phi(0)=\phi(i\bmod m)$. ∎
@@ Appendix C Regular Language Recognition Proof
[…]
@@
We now proceed by induction over $\ell$, defining the behavior of non-\$ tokens at layers that make up level $\ell$. The current group element $\delta^{\ell}_{i}$ stored at active token $i$ is, by inductive assumption, the cumulative product from $i-2^{\ell}+1$ to $i$. Let $\alpha^{\ell}_{i}$ denote that token $i$ is active. By Lemma˜7 we use a layer to store i1i-1 at token $i$. The next layer attends with query $\phi(i-1)$, key $\phi(j)$, and value $\delta^{\ell}_{j}$ to retrieve $\delta^{\ell}_{i-1}$, the group element stored at the previous token. Finally, another layer attends with query $\vec{1}$, key $\langle\phi(j)_{1},\alpha^{\ell}_{i}\rangle$, and value $\delta^{\ell}_{j-1}$ to retrieve the group element $\delta^{\ell}_{j^{*}}$ stored at the previous active token, which represents the cumulative product from $i-2\cdot 2^{\ell}+1$ to $i-2^{\ell}$. Next, we will use two layers to update $\delta^{\ell}_{i}\leftarrow\delta^{\ell+1}_{i}$ and $\delta^{\ell}_{j}\leftarrow\vec{0}$, which is achieved as follows. First, we assert there exists a single feedforward layer that uses a table lookup to compute $\delta^{\ell}_{j^{*}},\delta^{\ell}_{i}\mapsto d$ such that $d/\lVert d\rVert=\delta^{\ell}_{j^{*}}\cdot\delta^{\ell}_{i}=\delta^{\ell+1}_{i}$. Next, we invoke Lemma˜6 to construct a layer that adds $d$ to an empty cell of the residual stream and then another layer that deletes it. This second layer can now read both $\delta^{\ell}_{i},\delta^{\ell}_{j^{*}}$ and $\delta^{\ell+1}_{i}$ (from $d$) as input, and we modify it to add $\delta^{\ell+1}_{i}-\delta^{\ell}_{i}$ to $\delta^{\ell}_{i}$, changing its value to $\delta^{\ell+1}_{i}$. Similarly, we modify it to add $\delta^{\ell}_{j^{*}}$ to $\delta^{\ell}_{j^{*}}$ to set it to 0. A feedforward network then subtracts $\delta^{\ell}_{i}$ from the residual stream and adds $\delta^{\ell}_{i}\cdot\delta^{\ell}_{j}$. This requires at most 4 layers.
@@
To determine activeness in layer $\ell+1$, each token $i$ attends to its left to compute $c_{i}/i$, where $c_{i}$ is the prefix count of active tokens, inclusive of the current token. We then compute $\phi(c_{i}/i,1/i)=\phi(c_{i})$ and store $c_{i}$ it temporarily in the residual stream. At this point, we use Lemma˜1 to construct 7 layers that compute $c_{i}\bmod 2$ with no storage overhead. The current token is marked as active in layer $\ell+1$ iff $c_{i}=0\mod 2$, which is equivalent to checking whether $i=t\cdot 2^{\ell}-1$ for some $t$. In addition to updating the new activeness $\alpha^{\ell+1}_{i}$, we also persist store the previous activeness $\alpha^{\ell}_{i}$ in a separate cell of the residual stream and clear $c_{i}$. This requires at most 8 layers.
[…]
Finally, we justify the model size $m$ and feedforward width $w$. To represent a nondeterministic monoid element $Q\to Q$, which is defined by a matrix in $\{0,1\}^{\lvert Q\rvert\times\lvert Q\rvert}$, we must store $s_{\text{NFA}}=\lvert Q\rvert^{2}$ bits. For a deterministic monoid element, we can reduce this to $s_{\text{DFA}}=\lvert Q\rvert\log\lvert Q\rvert$ by storing the index of the unique $1$ in each row of the matrix. The model size is then $m=\mathrm{O}(s)$, which gives:
[…]
The feedforward network stores a lookup table enumerating all possible values of $s^{2}$ bits, returning $s$ bits in each case. This can be represented as long as $m=\mathrm{O}(s)$ and $w=\mathrm{O}(2^{s^{2}})$, which gives:
[…]
This finishes the proof. ∎